# On the MST-ratio: Theoretical Bounds and Complexity of Finding the Maximum

### v2

@@ 3 Maximum MST-ratio of Geometric Point Sets
On the other hand, considering a random coloring of $P$ with two colors, the probability for the edge $e_{i}$ being colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent. Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$. Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful. By the above argument and since there are $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
As $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, then:
$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|EMST(P)|-\frac{|EMST(P)|}{n-1},$
@@
Note that Theorem 4.3 holds also for general graphs.
The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true. The only instances we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-6\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 6$ points are sufficiently far from the core such that their pairwise distances are not less than their distances to the core. We believe those are the only point sets with an average EMST-ratio of less than $1$.
@@
As an example, we show that the average EMST-ratio for the point set in Figure 6 is less than $1$. The point set consists of $n-5$ points in a small circle through the origin whose radius is $\varepsilon\in{{o}}(\frac{1}{n2^{n}})$ (the core), along with $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin. In this example, clearly $|EMST(P)|\geq 5-{{o}}(n\varepsilon)$. In any bi-coloring of $P$ by red and blue if an outer point has the same color as one of the core points, then the edge connecting this outer point to a core point appears in one of the monochromatic EMSTs. Therefore, in most of the bi-colorings of $P$ where the core contains both colors, we have $|EMST(R)|+|EMST(B)|\leq 5+{{o}}(n\varepsilon)$. There are only $62$ many proper bi-colorings with a mono-chromatic core, and one can show that in each of them $|EMST(R)|+|EMST(B)|<5.1$. In fact, in $52$ of such colorings we even have $|EMST(R)|+|EMST(B)|<4.8$. Thus the average EMST-ratio for this point set is upper bounded by:
$\frac{(2^{n}-64)\times(5+{{o}}(n\varepsilon))+52\times(4.8)+10\times(5.1)}{(2^{n}-2)\times(5-{{o}}(n\varepsilon))}<1.$
Figure 6: A point set with average EMST-ratio smaller than $1$. The point set consists of $n-5$ points in a very small circle through the origin (the core), and $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin. The corresponding MST is depicted.
Question 4.5.
Can the point sets in ${{\mathbb{R}}}^{2}$ with an average EMST-ratio $<1$ be characterized?
5 Computational Experiments
@@
Figure 9, on the other hand, presents a comparison between a bipartite EMST-ratio computed using Prim’s algorithm and the average EMST-ratio obtained from $10^{5}$ experiments with point sets ranging from $5$ to $20$ points. Interestingly, in all computed experiments without exception, the computed bipartite EMST-ratio exceeds the average EMST-ratio. This suggests that bipartite colorings of a point set provide better EMST-ratios than what is expected from a random coloring. Although this is not always the case for finite point sets, we believe there are very few counterexamples.
Question 5.4.
@@
6 Discussion
In addition to the conjectures and questions posed in this paper, there are several other directions worth exploring further.
Coloring with Many Crossings.
For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$. Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring appears. In fact, for any point set, it is easy to find colorings with no crossings between the two MSTs. More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
Figure 10: An example of a point set and a bi-partition with chromatic crossing number $4$.
Question 6.1.
Question 6.2.
Conjecture 6.3.
Maximum MST Length over Subsets.
Related question to finding the maximum MST-ratio is the following. In the MAX-EMST-subset problem, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$. The question can be raised also in the abstract setting. Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST weight over all induced subgraphs $H\subseteq G$.
Question 6.4.
Are the problems MAX-MST-subset and MAX-EMST-subset NP-hard?
More Colors.
Similar questions can be asked for point sets with $c>2$ colors. To express the MST-ratio, we consider the total length of minimum spanning trees of points in each color divided by the length of a minimum spanning tree of the (uncolored) point set. Theorem 3.8 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space. For points in Euclidean space, this bound can be improved for $c\geq 4$. In particular, as $\rho_{d}\geq 0.577$ for every $d$, the maximum EMST-ratio is upper bounded by $\frac{c}{0.577}$. It remains unclear how the results in [9] on lattices can be generalized to more than $2$ colors. Additionally, exploring the average EMST-ratio for $c$ colors is an interesting direction to pursue.
Acknowledgements. The authors thank Ondřej Draganov for his valuable insights on the topic of average EMST-ratio.

### v3

@@ 3 Maximum MST-ratio of Geometric Point Sets
On the other hand, considering a random coloring of $P$ with two colors, the probability that the edge $e_{i}$ is colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent. Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$. Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful. By the above argument and since there are $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio $a$ is at least:
$a\geq|\text{EMST}(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
As we have $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, it can be concluded that $a\geq|\text{EMST}(P)|-\frac{1}{n-1}\cdot|\text{EMST}(P)|$, and hence the claim.
Note that Theorem 4.2 holds also for general graphs. The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true. The only instances we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-6\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 6$ points are sufficiently far from the core such that their pairwise distances are not less than their distances to the core. We believe these are the only point sets with an average EMST-ratio of less than $1$. As an example, in Section 0.A.2, we show that the average EMST-ratio for the point set in Figure 4.b is less than $1$.
Question B
Which point sets in ${{\mathbb{R}}}^{2}$ have an average EMST-ratio $<1$?
5 Discussion
In addition to the conjectures and questions posed in this paper, there are several other directions that are worth exploring further.
Coloring with Many Crossings.
For a point set $P$ in the plane, partitioned as $P=R\cup B$ of $P$, the Chromatic Crossing Number is the number of crossings between pairs of edges, one in $\text{EMST}(R)$ and one in $\text{EMST}(B)$. Intuitively, this also measures how mixed the coloring is; the higher the number, the more mingled it appears. In fact, for any point set, it is easy to find colorings with no crossings between the two MSTs. More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
Figure 5: A point set and a bi-partition with chromatic crossing number $4$.
Question C
Question D
Conjecture 3
Maximum MST Length over Subsets.
A related question to finding the maximum MST-ratio is the MAX-EMST-subset problem. Here, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$. This extends to the abstract setting as the MAX-MST-subset problem, where, given a weighted graph $G$, we seek the maximum MST weight over all induced subgraphs $H\subseteq G$.
Question E
Are MAX-MST-subset and MAX-EMST-subset NP-hard?
More Colors.
Similar questions can be asked for point sets with $c>2$ colors. To express the MST-ratio, we consider the total length of MST of points in each color divided by the length of an MST of the (uncolored) point set. Theorem 3.2 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space. For points in Euclidean space, this bound can be improved for $c\geq 4$. In particular, as $\rho_{d}\geq 0.577$ for every $d$, the maximum EMST-ratio is upper bounded by $\frac{c}{0.577}$. It remains unclear how the results in [10] on lattices can be generalized to more than $2$ colors. Furthermore, exploring the average EMST-ratio for $c$ colors is an interesting direction to pursue.
Appendix 0.A Appendix
0.A.1 Maximum EMST-ratio of Triangular Chain
In this section, we show that the maximum EMST-ratio for the triangular chain, shown in Figure 3 is $\frac{3k-2}{2k+1}$. Let $P$ be the point set of a triangular chain with $12k+$ 1 points and let us number the points of $P$ from left to right by $p_{1}$, $p_{2}$, $\ldots$, $p_{2k+1}$.
As $|\text{EMST}(P)|=2k$, to prove the claim we show that for any proper coloring of the chain to red and blue, we have $|\text{EMST}(R)|+|\text{EMST}(B)|\leq 3k-2$. We prove this by applying induction on $k$. The claim holds trivially for $k=1$, as any proper coloring has $|\text{EMST}(R)|+|\text{EMST}(B)|=1=3-2$.
Now, assume that the claim holds for $k$. Now we prove the claim for $k+1$. Assume we have a chain with $12(k+1)+$ 1 points. Let $B$ and $R$ be the set of blue points and red points in a proper coloring of $P$. We distinguish two cases:
a) If there exists $i\in\{1,3\}$, such that $p_{i}$, $p_{i+1}$ and $p_{i+2}$ have the same color: W.l.o.g. we assume these points belong to $R$. First, consider the case $i=1$. In this case, $R\setminus\{p_{1},p_{2}\}$ and $B$ is a proper coloring for $P\setminus\{p_{1},p_{2}\}$. Also, $T_{R}:=\text{EMST}(R\setminus\{p_{1},p_{2}\})\cup\{p_{1}p_{2},p_{2}p_{3}\}$ is a spanning tree of $R$. By induction hypothesis $|\text{EMST}(R\setminus\{p_{1},p_{2}\})|+|\text{EMST}(B)|\leq 3k-2$, which implies that:
$|\text{EMST}(R)|+|\text{EMST}(B)|\leq|T_{R}|+|\text{EMST}(B)|$
$\leq 3k-2+|p_{1}p_{2}|+|p_{2}p_{3}|=3k-1+2<3(k+1)-2,$
@@ 0.A.1 Maximum EMST-ratio of Triangular Chain
Now consider the case $i\neq 1$ (i.e. $i=3$). In this case, let $P^{\prime}$ be the triangular chain obtained from $P$ by removing $p_{1}$ and $p_{2}$. Let $R^{\prime}$ and $B^{\prime}$ be a proper coloring of $P^{\prime}$, where:
  • For every $j\geq 5$, $p_{j}\in R^{\prime}$ iff $p_{j}\in R$.
  • For every $j\in\{3,4\}$, $p_{j}\in R^{\prime}$ iff $p_{j-2}\in R$.
Note that $R^{\prime}$ and $B^{\prime}$ is a proper coloring of $P^{\prime}$ and hence by induction hypothesis:
$|\text{EMST}(R^{\prime})|+|\text{EMST}(B^{\prime})|\leq 3k-2.$
As $i\neq 1$, then $(p_{1}\cup p_{2})\cap B\neq\emptyset.$ We have $|\text{EMST}(R^{\prime})|=|\text{EMST}(R)|-2$ and $|\text{EMST}(B^{\prime})|+1\geq|\text{EMST}(B)|$. Therefore:
$|\text{EMST}(R)+|\text{EMST}(B)|\leq|\text{EMST}(R^{\prime})|+|\text{EMST}(B^{\prime})|+3\leq 3(k+1)-2.$
b) Otherwise: W.l.o.g. assume $p_{3}\in R$. First, assume $p_{1}$ and $p_{2}$ are in $B$. Since we are not in case a), then at least one of the points $p_{4}$ and $p_{5}$ is in $B$. Therefore, $R$ and $B\setminus\{p_{1},p_{2}\}$ is a proper coloring for $P\setminus\{p_{1},p_{2}\}$. Now $T_{B}:=\text{EMST}(B\setminus\{p_{1},p_{2}\})\cup\{p_{1}p_{2},p_{2}p_{i}\}$ is a spanning tree of $B$, where $p_{i}\in\{p_{4},p_{5}\}\cap B$. Note that by the induction hypothesis $|\text{EMST}(B\setminus\{p_{1},p_{2}\})|+|\text{EMST}(R)|\leq 3k-2$, therefore
$|\text{EMST}(R)|+|\text{EMST}(B)|\leq|T_{B}|+|\text{EMST}(R)|$
$\leq 3k-2+|p_{1}p_{2}|+|p_{2}p_{i}|=3k-1+|p_{2}p_{i}|<3(k+1)-2.$
Now assume that there exists $j\in\{1,2\}$ such that $p_{j}\in R$. Since we are not in case a), then $p_{3-j}$ is blue and at least one of the points $p_{4}$ and $p_{5}$ (say $p_{i}$) is in $B$. Therefore, $R$ and $B\setminus\{p_{1},p_{2}\}$ is a proper coloring for $P\setminus\{p_{1},p_{2}\}$. Now $T_{B}:=\text{EMST}(B\setminus\{p_{1},p_{2}\})\cup\{p_{j}p_{3},p_{3-j}p_{i}\}$ is a spanning tree of $B$. Note that by induction hypothesis $|\text{EMST}(B\setminus\{p_{1},p_{2}\})|+|\text{EMST}(R)|\leq 3k-2$, therefore
$|\text{EMST}(R)|+|\text{EMST}(B)|\leq|T_{B}|+|\text{EMST}(R)|$
$\leq 3k-2+|p_{j}p_{3}|+|p_{3-j}p_{i}|=3k-1+|p_{3-j}p_{i}|\leq 3(k+1)-2.$
0.A.2 A point set with average EMST-ratio $<1$
@@
In this section, we show that the average EMST-ratio for the point set described in Figure 4.b is smaller than $1$. The point set consists of $n-5$ points in a small circle through the origin whose radius is $\varepsilon\in{{o}}(\frac{1}{n2^{n}})$ (the core), along with $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin. Observe that each side of this pentagon is $s=2\sin{36}\approx 1.176$ and each diagonal is $d=\frac{1+\sqrt{5}}{2}s\approx 1.902$. In this example, clearly $|\text{EMST}(P)|\geq 5-{O}(n\varepsilon)$. In any bi-coloring of $P$ by red and blue if an outer point has the same color as one of the core points, then the edge connecting this outer point to a core point appears in one of the monochromatic EMSTs. Therefore, in $642^{n}-$ 64 bi-colorings of $P$ where the core contains both colors, we have $|\text{EMST}(R)|+|\text{EMST}(B)|\leq 5+{O}(n\varepsilon)$. There are only $622^{6}-2=$ 62 proper bi-colorings with a mono-chromatic core. Consider one of these $62$ colorings and assume the points in the core are red. In this case, if $50\leq n_{r}<$ 5 denotes the number of outer points that are red, then $|\text{EMST}(R)|=n_{r}+O(n\varepsilon)$. Also, if the outer blue points are consecutive vertices of the outer pentagon, $|\text{EMST}(B)|=(4-n_{r})s$ and otherwise $|\text{EMST}(B)|=(3-n_{r})s+d$. So in each of these colorings $|\text{EMST}(R)|+|\text{EMST}(B)|\leq d+s+2+{O}(n\varepsilon)<5.1$. Moreover, in $62-2\times 5=52$ of such colorings we even have $|\text{EMST}(R)|+|\text{EMST}(B)|\leq d+3+{O}(n\varepsilon)<4.91$. Thus, the average EMST-ratio for this point set is upper bounded by:
$\frac{(2^{n}-64)\cdot(5+{O}(n\varepsilon))+52\cdot(4.91)+10\cdot(5.1)}{(2^{n}-2)\cdot(5-{O}(n\varepsilon))}<1.$
0.A.3 Computational Experiments
@@
Figure 8, on the other hand, presents a comparison between a bipartite EMST-ratio computed using Prim’s algorithm and the average EMST-ratio obtained from $10^{5}$ experiments with point sets ranging from $5$ to $20$ points. Interestingly, in all computed experiments without exception, the computed bipartite EMST-ratio exceeds the average EMST-ratio. This suggests that bipartite colorings of a point set provide better EMST-ratios than what is expected from a random coloring. Although this is not always the case for finite point sets, we believe that there are very few counterexamples.
Question G